OpenStudy (anonymous):

Solve sin(x)(sinx + 1)=0.

OpenStudy (anonymous):

zero product property says that \(\sin x=0 \) or \(\sin x=-1 \) ....

OpenStudy (anonymous):

Where are you confused on what to do first?

OpenStudy (anonymous):

I'm not really sure... I don't know what to do first

OpenStudy (anonymous):

The advice that @mukushla provided is exactly how to go about doing this problem. Do you understand what he said?

OpenStudy (anonymous):

Yes, I understand the zero product property. So to solve that, what would I do? Would I have to do sin^-1(0) and sin^-1(-1) to find x?

OpenStudy (anonymous):

Oh no. So the first solution we have is \[\huge \sin(x)=0\] To solve for x we take the arcsin of both sides. Have you learned that yet?

OpenStudy (anonymous):

Yeah I learned it, but I don't really understand it

OpenStudy (anonymous):

s\[sinx=0 \implies x=\pm n \Pi \] where n is any integer

OpenStudy (anonymous):

If not, we should know for what x makes sin equal to zero. Hint: consult the unit circle. :D

OpenStudy (anonymous):

@MathFreak106 https://docs.google.com/open?id=0B98cXMBCs13BZThlZDAyMGUtZGIxMi00NjVhLWFjODktMGU4NjQ0ZDQzMWU5 This is a nice formula sheet i made a couple years ago.

OpenStudy (anonymous):

\[sinx=-1 \implies x=3\pi/2.....\]

OpenStudy (anonymous):

Lets look at arcsin for a moment. if sin(x)=y THEN, arcsin(y)=x Where x and y aren't coords but any number or function.

OpenStudy (anonymous):

Ok so arcsin(0)=x

OpenStudy (anonymous):

RIGHT! and the arcsin of a function will find you an angle. Do you see why?

OpenStudy (anonymous):

Kind of. How do you solve arcsin?

OpenStudy (anonymous):

Do you have a unit circle around you?

OpenStudy (anonymous):

Yep, from your worksheet :)

OpenStudy (anonymous):

great! Lets look at this example \[\huge \sin(x) = {1 \over 2}\] Can you find the two angles where that happens?

OpenStudy (anonymous):

Maybe it's better if i write it like: \[\huge \sin(\theta)= {1 \over 2}\]

OpenStudy (anonymous):

Umm 5pi/6?

OpenStudy (anonymous):

Ya, thats one of them. There's one more.

OpenStudy (anonymous):

I don't know what the other one is

OpenStudy (anonymous):

check out \[\pi \over 6\]

OpenStudy (anonymous):

Remember Theta can be in radians or degrees too. So looking at the unit circle again it could be 30 degrees or 150 degrees. So if sin(x)=0 that means it could be pi or 0, OR 180 degrees or 0 degrees. But if you look at the other side of my sheet you will see that arcsin has a range from \[\huge [-\pi/2, \pi/2]\] which means that it can ONLY be ONE of the choices we have.

OpenStudy (anonymous):